# Non-Abelian interference of topological edge states

### v1

@@ Abstract
Topological boundary states exhibit distinctive properties, including unidirectional propagation and noise robustness, which hold significant potential for advancing the performance of quantum science and technology. Here, we demonstrate the implementation of non-Abelian quantum interference and entanglement generation, protected by dual symmetries (time-independent inversion and time-dependent interchain), in coupled Su-Schrieffer-Heeger chains. Specifically, in a multi-chain system, we first achieve tunable topological transfer of a single particle, where the destination chain is selected by the braiding sequence. We then extend this to two particles, observing a non-Abelian Hong-Ou-Mandel interference that generates spatially entangled NOON states whose properties are dictated by the braiding order. Our work establishes an alternative pathway for exploring non-Abelian topology applied to quantum science and technology, enabled by the unique protection of time-dependent symmetry.
@@ I Introduction
In parallel to these developments, the role of fundamental quantum effects in such systems has garnered significant attention. Specifically, quantum entanglement and interference play a crucial role in the three major fields of quantum science and technology, including quantum communication, quantum computation, and quantum metrology [34, 35, 36]. Topological systems have been shown to facilitate the topological protection of quantum entanglement [37, 38, 39, 40], highlighting their potential for stabilizing quantum effects against external perturbations. For instance, the well-known Hong-Ou-Mandel (HOM) interference [41] of topological edge states has recently been observed in experiments on integrated photonic circuits [42]. Moreover, beyond experimental observations, several theoretical proposals have also been put forward to achieve interference of topological states, generate spatially entangled two-particle NOON states, and realize nonadiabatic topological transport protocols [43, 44, 45]. Notably, the frontier of this exploration is now extending into the realm of non-Abelian topology [46, 47, 48, 49, 50, 51]. In such systems, their non-commutative properties enriching the landscape of quantum phenomena. Despite this intriguing potential, research on quantum entanglement and interference mediated specifically by non-Abelian topological edge states remains largely unexplored.
@@
To address this gap, and motivated by recent advances in topological insulators with exotic gauge fields [52, 53, 54, 55] and quantum anomalous semimetals [56], we introduce a periodically modulated one-dimensional multi-chain system as a platform for non-Abelian quantum interference and entanglement generation via braiding of topological edge states. The non-Abelian characteristics of the platform are protected by a time-dependent interchain symmetry in two crucial aspects: First, it enables the formation of multiple edge-state branches with distinct symmetry eigenvalues, whose temporal braiding induces branch switching, which directly exhibits the non-Abelian properties. Second, it ensures the robust adiabatic evolution of states along a single branch, even when the energy gap to other branches closes transiently. Additionally, the dynamics are also governed by a time-independent inversion symmetry that protects the adiabatic evolution of parity-distinct edge states, with a relative phase accumulating to determine the arbitrary partition of particle transfer at the chain ends. Specifically, when each edge-state branch is initially localized on a separate chain (with uncoupled initial conditions), a particle injected into one end of a chain (as a superposition of parity-opposite states in its associated branch) can be transferred to the two ends of another chain at a tunable ratio (0 to 1). The non-Abelian character of this transport is evidenced by the fact that the choice of the recipient chain is determined by the sequence of braiding, while the final partition within that chain is controlled by the accumulated dynamical phase. When extended to two particles, our system allows for the observation of a non-Abelian HOM interference, a process in which the braiding sequence dictates the generated entangled state.
@@
This article is organized as follows. In Sec. II, we discuss the braiding of two pairs of topological edge states protected by the time-dependent interchain symmetry in a double-coupled SSH chain. In Sec. III, we extend our analysis to a triple-coupled SSH chain system and achieve topological transport, HOM interference, and entanglement generation based on the non-Abelian braiding of three pairs of edge states. Finally, a summary and discussion are provided in Sec. IV.
@@
II Braiding of two pairs of topological edge states in a double-coupled chain
@@
In this section, we focus on the braiding of two pairs of topological edge states in a double-chain system. While this system does not exhibit non-Abelian characteristics, its relative simplicity makes it an ideal starting point for introducing the fundamental mechanism. Specifically, we consider double-coupled Su-Schrieffer-Heeger (SSH) chains, labeled $A$ and $B$, with opposite global on-site energies. The system is governed by the Hamiltonian (with $\hbar=1$):
@@
where $\hat{\bm{b}}_{l}^{{\dagger}}=(\hat{b}_{Al}^{{\dagger}},\hat{b}_{Bl}^{{\dagger}})$ and $\hat{\bm{b}}_{l}=(\hat{b}_{Al},\hat{b}_{Bl})^{T}$ with $\hat{b}_{Al}^{{\dagger}}$ ($\hat{b}_{Bl}^{{\dagger}}$) and $\hat{b}_{Al}$ ($\hat{b}_{Bl}$) being the creation and annihilation operators acting on lattice site $l$ of SSH Chain AA ($B$), respectively. Furthermore, the intrachain hopping amplitudes are set as $v=v_{0}|\sin(\omega t)|$ and $w=1$ (taken as the energy unit), while the interchain hopping amplitude and the on-site energy offset are $J=\kappa\sin(2\omega t)$ and $\Delta=\kappa[1+\cos(2\omega t)]$, respectively. With these parameters, this time-dependent Hamiltonian has a period $T=\pi/\omega$. An illustrative plot for this double-chain system with $2L$ lattice sites is provided in Fig. 1.
@@
Both $\bm{P}$ and $\bm{S}_{D}(t)$ possess two eigenvalues: $\bm{P}$ has $\pm 1$ (corresponding to odd/even parity), and $\bm{S}_{D}(t)$ has $S_{D,\text{I}}(t)=-S_{D,\text{II}}(t)=e^{-i\omega t}$. We note that the eigenvalues of $\bm{S}_{D}(t)$ exhibit a swapping property: $S_{D,\text{I}/\text{II}}(T)=S_{D,\text{II}/\text{I}}(0)$, with their associated eigenstates behaving identically. This induces intriguing phenomena in the system’s static energy spectrum and dynamical evolution, as discussed below.
II.1 Twisted energy spectrum and braiding of two pairs of edge states
Figure 2: Braiding of two pairs of topological edge states. (a) Energy spectrum of the double-coupled SSH chains. Black solid lines represent the bulk states, and colored lines represent pairs of edge states. (b) Two branches of the expectation value $\langle\bm{S}_{D}\rangle$. (c) Abelian braiding processes of two pairs of edge states in the ($t$, $\text{Im}[\langle\bm{S}_{D}\rangle]$, $E$) space, corresponding to the braiding operators $\tau^{-1}_{1}$. In all panels, the color denotes the argument of the expectation value $\langle\bm{S}_{D}\rangle$. The parameters set as $L=14$, $w=1$, $v_{0}=0.5$, and $\kappa=0.2$.
@@ II.1 Twisted energy spectrum and braiding of two pairs of edge states
More intuitively, Fig. 2(b) visualizes the real and imaginary parts of the edge states’ expectation value with respect to $\bm{S}_{D}(t)$ (denoted as $\langle\bm{S}_{D}\rangle=\langle\psi|\bm{S}_{D}(t)|\psi\rangle$), and Fig. 2(c) depicts the variation of the branches formed by the imaginary parts of $\langle\bm{S}_{D}\rangle$ and the energies. Here, the braiding behavior of the edge states is clearly visible. Specifically, within the single period depicted in Fig. 2(c), the braid operation $\tau^{-1}$ occurs, where we define $\tau_{n}$ ($\tau_{n}^{-1}$) as the braid operation in which the $n$-th strand crosses under (above) the $(n+1)$-th strand. This induces the aforementioned switching of the edge states’ energy and their association with $\bm{S}_{D}(t)$.
Given the periodicity of the Hamiltonian and the symmetry operator $\bm{S}_{D}(t)$, i.e., $\hat{H}_{D}(T)=\hat{H}_{D}(0)$ and $\bm{S}_{D}(T)=\bm{S}_{D}(0)$, we ultimately conclude that the two pairs of edge states swap after one period, following $|\psi_{\pm,\text{I}/\text{II}}(T)\rangle=|\psi_{\pm,\text{II}/\text{I}}(0)\rangle$. This phenomenon is more pronounced under the parameter variation considered herein, which has been presented below Eq. (1): at the initial and final moments of the period, the intrachain SSH segment of the system is in the fully dimerized limit ($v=0$), and the interchain coupling is decoupled ($J=0$). Such a configuration implies that at these two instants, the eigenedge states are fully localized odd-even superposition states at the two ends of a single chain. By further incorporating information on the energy and the expectation values of the symmetry operator, we obtain the following expressions:
@@ II.2 Symmetry-protected adiabatic evolution
We now turn to the adiabatic evolution of the edge states under the protecting of inversion symmetry and time-dependent interchain symmetry. Firstly, the parity of the states is conserved despite their extremely small energy difference during dynamical evolution, owing to inversion symmetry. Similarly, the interchain symmetry can also protect the adiabatic evolution of the edge states. This protection is manifested in the independent evolution of edge states in the two branches. Even though these two branches exhibit energy crossing, this feature would imply infinite adiabatic velocity in the absence of symmetry protection. We note that, in contrast to time-independent symmetries (such as inversion symmetry), this adiabatic evolution under the protection of time-dependent symmetry still requires a sufficient long (yet not infinite) evolution time, as discussed in the Appendix. Thereby, by combining them with the braiding properties (another effect brought by the time-dependent interchain symmetry), the edge states can propagate from the left end to the right end of the system during adiabatic evolution, and also transition from one chain to the other.
@@
III Non-Abelian braiding and Hong-Ou-Mandel interference in a triple-coupled chain
@@
The braiding of two pairs of topological edge states in Sec. II demonstrates the protection of time-dependent interchain symmetry, which we now extend to the case of three pairs described by non-Abelian braiding. We consider triple-coupled SSH chains denoted as $A$, $B$, and $C$, which are described by the following Hamiltonian (with $\hbar=1$):
@@
The eigenvalues $\pm 1$ of $\bm{P}$ can be used to denote the even and odd parity of the edge states, respectively. What’s more, the eigenvalues $S_{T,\text{I}}(t)=e^{2i\pi/3}S_{T,\text{II}}(t)=e^{-2i\pi/3}S_{T,\text{III}}(t)=e^{-2i\omega t/3}$ of $\bm{S}_{T}(t)$ exhibit a cyclic swapping property: $S_{T,\text{I}/\text{II}/\text{III}}(T)=S_{T,\text{II}/\text{III}/\text{I}}(0)$. This property induces non-Abelian braiding of pairs of edge states, as discussed below.
III.1 Non-Abelian braiding of three pairs of edge states
@@
Figure 5: Non-Abelian braiding of three pairs of edge states. (a) Energy spectrum of edge states for the triplet-coupled SSH chains. (b) Two branches of the expectation value $\langle\bm{S}_{T}\rangle$. (c), (d) Non-Abelian braiding processes of three pairs of edge states in the ($t$, $\text{Im}[\langle\bm{S}_{T}\rangle]$, $E$) and ($t$, $\text{Im}[\langle\bm{S^{*}}_{T}\rangle]$, $E$) space, corresponding to the braiding operators $\tau^{-1}_{1}\tau^{-1}_{2}$ (panel (c)) and $\tau^{-1}_{2}\tau^{-1}_{1}$ (panel (d)). In all panels, the color indicates the argument of the expectation value $\langle\bm{S}_{T}\rangle$ or $\langle\bm{S^{*}}_{T}\rangle$. The parameters set as $L=14$, $w=1$, $v_{0}=0.5$, and $\eta=0.2$.
What’s more, during a single period, the variation of the branches (which are formed by the imaginary parts of $\langle\bm{S}_{T}\rangle$ and energy) is described by the non-Abelian braid operation $\tau_{1}^{-1}\tau_{2}^{-1}$, as depicted in Fig. 5(c). Leveraging the periodicity of both the Hamiltonian and the symmetry operator $\bm{S}_{T}(t)$, i.e., $\hat{H}_{T}(T)=\hat{H}_{T}(0)$ and $\bm{S}_{T}(T)=\bm{S}_{T}(0)$, we ultimately conclude that the three pairs of edge states undergo cyclic mutual transformation over one period, adhering to the relation $|\Psi_{\pm,\text{I}/\text{II}/\text{III}}(T)\rangle=|\Psi_{\pm,\text{II}/\text{III}/\text{I}}(0)\rangle$. To better exhibit the cyclic mutual swapping, we choose the parameters such that at both the initial and final moments of the period, the intrachain SSH segments of the system are in the fully dimerized limit ($v=0$), and the interchain couplings are decoupled ($J_{AB}=J_{AC}=J_{BC}=0$, see Eq. (42)). The edge states at these two instants are fully localized as odd-even superposition states at the two ends of individual chains, as specified below:
@@
The quintessential effect of non-Abelian braiding manifests as the operation-sequence-dependent swapping of edge states. To this end, we perform a time reversal operation on the Hamiltonian $\hat{H}_{T}(t)$ and the symmetry operator $\hat{\mathcal{S}}_{T}(t)$, that is
@@
As a result, during a single period, the variation of the branches exhibits a distinct pattern. Specifically, for branches formed by the imaginary parts of $\langle\bm{S}^{*}_{T}\rangle$ and energy, this variation is described by the non-Abelian braid operation $\tau_{2}^{-1}\tau_{1}^{-1}$, as depicted in Fig. 5(d). The cyclic mutual swapping of the edge states is distinctly different from that described by $\tau_{1}^{-1}\tau_{2}^{-1}$ (see Eq. (48)) and is given by
@@
Obviously, after traversing a period, starting with the same initial state, the final state governed by $\tau_{2}^{-1}\tau_{1}^{-1}$ differs from that governed by $\tau_{1}^{-1}\tau_{2}^{-1}$. This is precisely the manifestation of non-Abelian braiding. In this subsection, we have discussed symmetry-protected non-Abelian braiding; below, we will demonstrate non-Abelian topological transport based on this non-Abelian braiding.

### v2

@@ Abstract
Topological boundary states exhibit distinctive properties, including unidirectional propagation and noise robustness, which hold significant potential for advancing the performance of quantum science and technology. Here, we demonstrate the implementation of non-Abelian quantum interference and entanglement generation, protected by dual symmetries (time-independent inversion and time-dependent interchain), in coupled Su-Schrieffer-Heeger chains. Specifically, in a multi-chain system, we first achieve tunable topological transfer of a single particle, where the destination chain is selected by the permutation sequence. We then extend this to two particles, observing a non-Abelian Hong-Ou-Mandel interference that generates spatially entangled NOON states whose properties are dictated by the permutation sequence. Our work establishes an alternative pathway for exploring non-Abelian topology applied to quantum science and technology, enabled by the unique protection of time-dependent symmetry.
@@ I Introduction
In parallel to these developments, the role of fundamental quantum effects in such systems has garnered significant attention. Specifically, quantum entanglement and interference play a crucial role in the three major fields of quantum science and technology, including quantum communication, quantum computation, and quantum metrology [34, 35, 36]. Topological systems have been shown to facilitate the topological protection of quantum entanglement [37, 38, 39, 40], highlighting their potential for stabilizing quantum effects against external perturbations. For instance, the well-known Hong-Ou-Mandel (HOM) interference [41] of topological edge states has recently been observed in experiments on integrated photonic circuits [42]. Moreover, beyond experimental observations, several theoretical proposals have also been put forward to achieve interference of topological states, generate spatially entangled two-particle NOON states, and realize nonadiabatic topological transport protocols [43, 44, 45]. Notably, the frontier of this exploration is now extending into the realm of non-Abelian topology [46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61]. In such systems, their non-commutative properties enriching the landscape of quantum phenomena. Despite this intriguing potential, research on quantum entanglement and interference mediated specifically by non-Abelian topological edge states remains largely unexplored.
@@
To address this gap, and motivated by recent advances in topological insulators with exotic gauge fields [62, 63, 64, 65] and quantum anomalous semimetals [66], we introduce a periodically modulated one-dimensional multi-chain system as a platform for non-Abelian quantum interference and entanglement generation via permutation of topological edge states. The non-Abelian characteristics of the platform are protected by a time-dependent interchain symmetry in two crucial aspects: First, it enables the formation of multiple edge-state branches with distinct symmetry eigenvalues, whose temporal permutation induces branch switching, which directly exhibits the non-Abelian properties. Second, it ensures the robust adiabatic evolution of states along a single branch, even when the energy gap to other branches closes transiently. Additionally, the dynamics are also governed by a time-independent inversion symmetry that protects the adiabatic evolution of parity-distinct edge states, with a relative phase accumulating to determine the arbitrary partition of particle transfer at the chain ends. Specifically, when each edge-state branch is initially localized on a separate chain (with uncoupled initial conditions), a particle injected into one end of a chain (as a superposition of parity-opposite states in its associated branch) can be transferred to the two ends of another chain at a tunable ratio (0 to 1). The non-Abelian character of this transport is evidenced by the fact that the choice of the recipient chain is determined by the sequence of permutation, while the final partition within that chain is controlled by the accumulated dynamical phase. When extended to two particles, our system allows for the observation of a non-Abelian HOM interference, a process in which the permutation sequence dictates the generated entangled state.
@@
This article is organized as follows. In Sec. II, we discuss the permutation of two pairs of topological edge states protected by the time-dependent interchain symmetry in a double-coupled SSH chain. In Sec. III, we extend our analysis to a triple-coupled SSH chain system and achieve topological transport, HOM interference, and entanglement generation based on the non-Abelian permutation of three pairs of edge states. Finally, a summary and discussion are provided in Sec. IV.
@@
II Permutation of two pairs of topological edge states in a double-coupled chain
@@
In this section, we focus on the permutation of two pairs of topological edge states in a double-chain system. While this system does not exhibit non-Abelian characteristics, its relative simplicity makes it an ideal starting point for introducing the fundamental mechanism. Specifically, we consider double-coupled Su-Schrieffer-Heeger (SSH) chains, labeled $A$ and $B$, with opposite global on-site energies. The system is governed by the Hamiltonian (with $\hbar=1$):
@@
where $\hat{\bm{b}}_{l}^{{\dagger}}=(\hat{b}_{Al}^{{\dagger}},\hat{b}_{Bl}^{{\dagger}})$ and $\hat{\bm{b}}_{l}=(\hat{b}_{Al},\hat{b}_{Bl})^{T}$ with $\hat{b}_{Al}^{{\dagger}}$ ($\hat{b}_{Bl}^{{\dagger}}$) and $\hat{b}_{Al}$ ($\hat{b}_{Bl}$) being the creation and annihilation operators acting on lattice site $l$ of SSH Chain AA ($B$), respectively. Furthermore, the intrachain hopping amplitudes are set as $v=v_{0}|\sin(\omega t)|$ and $w=1$ (taken as the energy unit), while the interchain hopping amplitude and the on-site energy offset are $J=\kappa\sin(2\omega t)$ and $\Delta=\kappa[1+\cos(2\omega t)]$, respectively. In this work, we focus on the topological phase with $|v|<|w|$, where topological edge states are localized at the two ends of the chains, with a decay index of $(v/w)^{2}$ [44]. With these parameters, this time-dependent Hamiltonian has a period $T=\pi/\omega$. An illustrative plot for this double-chain system with $2L$ lattice sites is provided in Fig. 1.
@@
Both $\bm{P}$ and $\bm{S}_{D}(t)$ possess two eigenvalues: $\bm{P}$ has $\pm 1$ (corresponding to odd/even parity), and $\bm{S}_{D}(t)$ has $S_{D,\text{I}}(t)=-S_{D,\text{II}}(t)=e^{-i\omega t}$. We note that the eigenvalues of $\bm{S}_{D}(t)$ exhibit a swapping property: $S_{D,\text{I}/\text{II}}(T)=S_{D,\text{II}/\text{I}}(0)$, with their associated eigenstates behaving identically. This induces intriguing phenomena in the system’s static energy spectrum and dynamical evolution, as discussed below. It is worth noting that when constructing the double-coupled SSH chains, we first adopt a well-studied non-Hermitian Hamiltonian of a two-level system with states featuring braiding characteristics [67, 68, 69, 70] as the symmetry operator of our system. We then reverse-engineer the interchain term $h_{2}$ based on the commutation relation, from which the Hamiltonian $\hat{H}_{D}$ (presented in Eq. (1)) is derived. Notably, this method is equally applicable to the triple-coupled SSH chains discussed subsequently.
II.1 Twisted energy spectrum and permutation of two pairs of edge states
Figure 2: permutation of two pairs of topological edge states. (a) Energy spectrum of the double-coupled SSH chains. Black solid lines represent the bulk states, and colored lines represent pairs of edge states. (b) Two branches of the expectation value $\langle\bm{S}_{D}\rangle$. (c) Abelian braiding processes of two pairs of edge states in the ($t$, $\text{Im}[\langle\bm{S}_{D}\rangle]$, $E$) space, corresponding to the braiding operators $\tau^{-1}_{1}$. In all panels, the color denotes the argument of the expectation value $\langle\bm{S}_{D}\rangle$. The parameters set as $L=14$, $w=1$, $v_{0}=0.5$, and $\kappa=0.2$.
@@ II.1 Twisted energy spectrum and permutation of two pairs of edge states
More intuitively, Fig. 2(b) visualizes the real and imaginary parts of the edge states’ expectation value with respect to $\bm{S}_{D}(t)$ (denoted as $\langle\bm{S}_{D}\rangle=\langle\psi|\bm{S}_{D}(t)|\psi\rangle$), and Fig. 2(c) depicts the variation of the branches formed by the imaginary parts of $\langle\bm{S}_{D}\rangle$ and the energies. Here, the braiding behavior of the edge states is clearly visible. Specifically, within the single period depicted in Fig. 2(c), the braid operation $\tau_{1}^{-1}$ occurs, where we define $\tau_{n}$ ($\tau_{n}^{-1}$) as the braid operation in which the $n$-th strand crosses under (above) the $(n+1)$-th strand. This induces the aforementioned switching of the edge states’ energy and their association with $\bm{S}_{D}(t)$. The braiding processes in the ($t$, $\text{Im}[\langle\bm{S}_{D}\rangle]$, $E$) parameter space result in the swap of two pairs of topological edge states, as discussed in the following. It is worth noting that the phase factor in Eq. (10) can be taken as $\mathrm{e}^{\mathrm{i}(2n+1)\omega t}$ with $n\in\mathbb{Z}$ under the condition that the time-dependent interchain symmetry operator $\hat{\mathcal{S}}_{D}(t)$ and the Hamiltonian $\hat{H}_{D}$ share the same period, and different values of $n$ may reverse the braiding operator (i.e., $\tau_{1}\leftrightarrow\tau_{1}^{-1}$). However, both $\tau_{1}$ and $\tau_{1}^{-1}$ can be mapped to the permutation operation $\mathcal{M}_{1}$, where $\mathcal{M}_{n}$ is defined as the permutation operation that swaps the $n$-th pair of topological edge states with the $(n+1)$-th pair. It is convenient to describe the swap of these two pairs of edge states using the permutation operation.
Given the periodicity of the Hamiltonian and the symmetry operator $\bm{S}_{D}(t)$, i.e., $\hat{H}_{D}(T)=\hat{H}_{D}(0)$ and $\bm{S}_{D}(T)=\bm{S}_{D}(0)$, we ultimately conclude that the two pairs of edge states swap after one period, following $|\psi_{\pm,\text{I}/\text{II}}(T)\rangle=|\psi_{\pm,\text{II}/\text{I}}(0)\rangle$ described by the permutation operation $\mathcal{M}_{1}$. This phenomenon is more pronounced under the parameter variation considered herein, which has been presented below Eq. (1): at the initial and final moments of the period, the intrachain SSH segment of the system is in the fully dimerized limit ($v=0$), and the interchain coupling is decoupled ($J=0$). Such a configuration implies that at these two instants, the eigenedge states are fully localized odd-even superposition states at the two ends of a single chain. By further incorporating information on the energy and the expectation values of the symmetry operator, we obtain the following expressions:
@@ II.2 Symmetry-protected adiabatic evolution
We now turn to the adiabatic evolution of the edge states under the protecting of inversion symmetry and time-dependent interchain symmetry. Firstly, the parity of the states is conserved despite their extremely small energy difference during dynamical evolution, owing to inversion symmetry. Similarly, the interchain symmetry can also protect the adiabatic evolution of the edge states. This protection is manifested in the independent evolution of edge states in the two branches. Even though these two branches exhibit energy crossing, this feature would imply infinite adiabatic velocity in the absence of symmetry protection. We note that, in contrast to time-independent symmetries (such as inversion symmetry), this adiabatic evolution under the protection of time-dependent symmetry still requires a sufficient long (yet not infinite) evolution time, as discussed in the Appendix. Thereby, by combining them with the permutation properties (another effect brought by the time-dependent interchain symmetry), the edge states can propagate from the left end to the right end of the system during adiabatic evolution, and also transition from one chain to the other.
@@
III Non-Abelian permutation and Hong-Ou-Mandel interference in a triple-coupled chain
@@
The permutation of two pairs of topological edge states in Sec. II demonstrates the protection of time-dependent interchain symmetry, which we now extend to the case of three pairs described by non-Abelian permutation. We consider triple-coupled SSH chains denoted as $A$, $B$, and $C$, which are described by the following Hamiltonian (with $\hbar=1$):
@@
The eigenvalues $\pm 1$ of $\bm{P}$ can be used to denote the even and odd parity of the edge states, respectively. What’s more, the eigenvalues $S_{T,\text{I}}(t)=e^{2i\pi/3}S_{T,\text{II}}(t)=e^{-2i\pi/3}S_{T,\text{III}}(t)=e^{-2i\omega t/3}$ of $\bm{S}_{T}(t)$ exhibit a cyclic swapping property: $S_{T,\text{I}/\text{II}/\text{III}}(T)=S_{T,\text{II}/\text{III}/\text{I}}(0)$. This property induces non-Abelian permutation of pairs of edge states, as discussed below.
III.1 Non-Abelian permutation of three pairs of edge states
@@
Figure 5: Non-Abelian permutation of three pairs of edge states. (a) Energy spectrum of edge states for the triplet-coupled SSH chains. (b) Two branches of the expectation value $\langle\bm{S}_{T}\rangle$. (c), (d) Non-Abelian braiding processes of three pairs of edge states in the ($t$, $\text{Im}[\langle\bm{S}_{T}\rangle]$, $E$) and ($t$, $\text{Im}[\langle\bm{S^{*}}_{T}\rangle]$, $E$) space, corresponding to the braiding operators $\tau^{-1}_{1}\tau^{-1}_{2}$ (panel (c)) and $\tau^{-1}_{2}\tau^{-1}_{1}$ (panel (d)). In all panels, the color indicates the argument of the expectation value $\langle\bm{S}_{T}\rangle$ or $\langle\bm{S^{*}}_{T}\rangle$. The parameters set as $L=14$, $w=1$, $v_{0}=0.5$, and $\eta=0.2$.
What’s more, during a single period, the variation of the branches (which are formed by the imaginary parts of $\langle\bm{S}_{T}\rangle$ and energy) is described by the non-Abelian braid operation $\tau_{1}^{-1}\tau_{2}^{-1}$, as depicted in Fig. 5(c). Leveraging the periodicity of both the Hamiltonian and the symmetry operator $\bm{S}_{T}(t)$, i.e., $\hat{H}_{T}(T)=\hat{H}_{T}(0)$ and $\bm{S}_{T}(T)=\bm{S}_{T}(0)$, we ultimately conclude that the three pairs of edge states undergo cyclic mutual transformation over one period, adhering to the relation $|\Psi_{\pm,\text{I}/\text{II}/\text{III}}(T)\rangle=|\Psi_{\pm,\text{II}/\text{III}/\text{I}}(0)\rangle$ described by the non-Abelian permutation operation $\mathcal{M}_{1}\mathcal{M}_{2}$. To better exhibit the cyclic mutual swapping, we choose the parameters such that at both the initial and final moments of the period, the intrachain SSH segments of the system are in the fully dimerized limit ($v=0$), and the interchain couplings are decoupled ($J_{AB}=J_{AC}=J_{BC}=0$, see Eq. (42)). The edge states at these two instants are fully localized as odd-even superposition states at the two ends of individual chains, as specified below:
@@
The quintessential effect of non-Abelian permutation manifests as the operation-sequence-dependent swapping of edge states. To this end, we perform a time reversal operation on the Hamiltonian $\hat{H}_{T}(t)$ and the symmetry operator $\hat{\mathcal{S}}_{T}(t)$, that is
@@
As a result, during a single period, the variation of the branches exhibits a distinct pattern. Specifically, for branches formed by the imaginary parts of $\langle\bm{S}^{*}_{T}\rangle$ and energy, this variation is described by the non-Abelian braid operation $\tau_{2}^{-1}\tau_{1}^{-1}$, as depicted in Fig. 5(d). The cyclic mutual swapping of the edge states is distinctly different from that described by $\mathcal{M}_{1}\mathcal{M}_{2}$ (see Eq. (48)) and is given by
@@
Obviously, after traversing a period, starting with the same initial state, the final state governed by $\mathcal{M}_{2}\mathcal{M}_{1}$ differs from that governed by $\mathcal{M}_{1}\mathcal{M}_{2}$. This is precisely the manifestation of non-Abelian permutation. In this subsection, we have discussed symmetry-protected non-Abelian permutation; below, we will demonstrate non-Abelian topological transport based on this non-Abelian permutation.